OpenStudy (anonymous):

how much would $500 invested at 4% interest compounded monthly be worth after 5 years

OpenStudy (anonymous):

The monthly interest rate is given by \(4\%/12\). The number of months it compounds is \(5\times 12\).

OpenStudy (anonymous):

The normal compounding interest formula is: \[ p(t) = p_0(1+r)^t \]But in this case, we make that adjustment: \[ p(t) = p_0\left(1+\frac rn\right)^{nt} \]Where \(n=12\) because there are \(12\) months in a year.

OpenStudy (anonymous):

And \(r=4\% = 0.04\), and \(p_0=\$500\).

OpenStudy (anonymous):

They want \(t=5\), because 5 years.

OpenStudy (anonymous):

\[ p(5) = 500\left(1+\frac{0.04}{12}\right)^{12(5)} \]